OpenStudy (anonymous):

Which of the following ratios is not equivalent to 6:10? 3/5 9/15 48/80 24/45

OpenStudy (anonymous):

Do you know what are "equivalent ratios" ?

OpenStudy (anonymous):

its 24/45 right

OpenStudy (anonymous):

its last one for sure

OpenStudy (anonymous):

Good work @lawson4238

OpenStudy (anonymous):

as 6 : 10 = 6/10 = 6*4/10*4 = 24/40 So it must be 24/40 and NOT 24/45 and hence you are right with the option D)

OpenStudy (anonymous):

ye so u jus reduce to the lowest terms to find out if it equal

OpenStudy (anonymous):

Yes like this : \[\large{\frac{24}{45} = \frac{3*8}{3*15} = \frac{\cancel{3}^1 * 8 }{\cancel{3}^1 * 15} = \frac{8}{15} = \textbf{HCF of 8 and 15 is 1} }\] Therefore 8/15 is the most simplified form of 24/45 and therefore, 24/45 is not equal to 6/10